0.25 mole of hydrogen chloride was dissolved in distilled water and the volume made up of 0.50 dm3. If 15.0 cm3 of the solution requires 12.50 cm3 of aqueous sodium trioxocarbonate (IV) for neutralization, calculate the concentration of the alkaline solution
A. 0.30 mol dm3
B. 0.40 mol dm3
C. 0.50 mol dm3
D. 0.60 mol dm3

Please show working....
1
Expert's answer
2020-04-19T07:10:03-0400

Na2CO3 + 2HCl -->2NaCl + H2O+CO2;

The quantity of HCl that reacted is: n(HCl)=15*0.25/50=0.075 mol

The quantity of Na2CO3 that reacted is: n(Na2CO3)=1/2*n(HCl)=0.0375 mol;

C(Na2CO3)=n(Na2CO3)/V(solution)=0.0375/0.0125= 3 mol/dm3